Official Report Week Ending March 11, 2007

This forum is dedicated to the preservation of the WWII submarine USS Torsk in Baltimore, MD

Moderators: Gilbert, Bruce

Official Report Week Ending March 11, 2007

Postby Dave T » Tue Mar 13, 2007 7:18 pm

Summary of shipboard & offsite work for the week ending March 11, 2007

Crew on board Saturday: Frank, Dave, Chief Mike, Chief Bill, Big Al, Tim, Gilbert, Cathy, Larry, Chuck, Clayton, & Rich Pekelney from Pampanito in San Francisco.

Rich Pekelney visited the boat while he was on the east coast for a conference. Rich brought along hundreds of excellent photos of Pampanito’s recent drydocking and yard upkeep period as well as a lot of before and after shots of hull and superstructure components. It made us even more aware of the need for Torsk to have a yard period as soon as possible. Ship’s work ground to a halt as Rich ran some excellent videos on his laptop computer of live action on board ex USS Tang and ex USS Cutlass, both now sailing under foreign flags (Tang-Turkey & Cutlass-Taiwan) They both look great. Shots of surfaces, dives, and snorkeling taken from inside and outside the boats held everyone’s attention and brought back a lot of memories. Rich personally hired a professional film crew for some of the footage and has captured interesting and important parts of our submarine heritage. We all grabbed our ears with the boat crew each time the head valve plunged beneath a heavy sea. Thanks for sharing with us Rich.

Some boat work was actually accomplished:

Cathy, Frank, Chuck, and Dave worked on the Maneuvering restoration project. Dave obtained paint color sample chips during the week and the Maneuvering Project Team made selections for various equipment and panels based on historic photos and 1960s crew memories. The basic color theme is Navy Gray. There’s a surprise.

Gilbert did general housekeeping and helped haul trash out of LLAB for disposal.

Big Al relamped the boat as necessary, worked on the Emergency Lighting system switches, and obtained a number of potential 1MC replacement transformers during the week.

Larry worked on WT doors, did general housekeeping and ran an air charge.

Chief Mike and Clayton disassembled and disposed of the large power transformer we had stowed in LLAB. We were happy several years ago when we got it down the hatch in to LLAB. Now, we’re just as happy to have it gone. They also continued cleaning and organizing in LLAB. Gilbert and Big Al helped haul the trash topside.

Frank & Dave took more measurements for the FTR ductwork needed to install the new AC system for FB and FTR. Installation of the system is expected to begin in May. Hopefully, we’ll have cool air flowing into FB before summer begins.

Tim and Joe continued restoration of the AER gaugeboards. Pretty soon they’ll have AER looking as good as FER.

Chief Bill continued working in Plot on the vacuum tube inventory with testing and sorting of tubes.

*****************
Onsite Volunteer Hours - 87
Offsite Volunteer Hours - 9
Total Week Hours - 96
Year To Date Volunteer Hours - 612

That’s it for the week,
Dave Thames
Dave T
 
Posts: 88
Joined: Fri Dec 01, 2006 4:43 pm
Location: Bawlmer

Return to General Discussion

Who is online

Users browsing this forum: Al F, Baidu [Spider] and 5 guests